OpenStudy (anonymous):

Simplify: 2 + 6x - y - 2x + 8 + 4y A. 4x + 3y + 10 B. 8x + 4y + 10 C. 4x - 5y +10 D. -8x + 5y + 10

OpenStudy (ivettef365):

just combine like terms you have 6x - 2x = ? 4y - y = ? 2 + 8 = ?? those are the terms that you can combine

OpenStudy (anonymous):

its A. 4x + 3y + 10